# Adaptive Incentive Design with Learning Agents

### v2

@@ 2.1.2 Non-atomic Games
Consider a game $\tilde{G}$ with a finite set of player populations $\tilde{\mathcal{I}}$. Each population $i\in\tilde{\mathcal{I}}$ is comprised of a continuum set of (infinitesimal) players with mass $\tilde{M}_{i}>0$. Every (infinitesimal) player in any population can choose an action in a finite set $\tilde{S}_{i}$. The strategy distribution of population $i\in\tilde{\mathcal{I}}$ is $\tilde{x}_{i}=(\tilde{x}_{i}^{j})_{j\in\tilde{S}_{i}}$, where $\tilde{x}_{i}^{j}$ is the mass of players in population $i$ who choose action $j\in\tilde{S}_{i}$. Then, the set of all strategy distributions of population $i$ is $\tilde{X}_{i}=\left\{\tilde{x}_{i}\in\mathbb{R}^{|\tilde{S}_{i}|}|\sum_{j\in\tilde{S}_{i}}\tilde{x}_{i}^{j}=\tilde{M}_{i},~{}\tilde{x}_{i}^{j}\geqslant 0,\forall j\in\tilde{S}_{i}\right\}$. The strategy distribution of all populations is $\tilde{x}=(\tilde{x}_{i})_{i\in\tilde{\mathcal{I}}}\in\tilde{X}=\prod_{i\in\tilde{\mathcal{I}}}\tilde{X}_{i}$. We define $\tilde{S}=\prod_{i\in\tilde{\mathcal{I}}}\tilde{S}_{i}$. Given a strategy distribution $\tilde{x}\in\tilde{X}$, the cost of players in population $i\in\tilde{\mathcal{I}}$ for choosing action $j\in\tilde{S}_{i}$ is ~$\tilde{\ell}_{i}^{j}(\tilde{x})$. We denote ~$\tilde{\ell}_{i}(\tilde{x})=(\tilde{\ell}_{i}^{j}(\tilde{x}))_{j\in\tilde{S}_{i}}$ as the vector of costs for each $i\in\tilde{\mathcal{I}}$.
A social planner designs incentives by setting a payment $\tilde{p}_{i}^{j}$ for players in population $i$ who choose action $j\in\tilde{S}_{i}$. Consequently, given the incentive mechanism $\tilde{p}=({\tilde{p}_{i}^{j}})_{j\in S_{i},i\in\tilde{\mathcal{I}}}$, the total cost experience by any player, in population $i\in\tilde{\mathcal{I}}$, that chooses action $j\in\tilde{S}_{i}$ is
@@
A strategy distribution ${\tilde{x}^{\ast}}(\tilde{p})\in\tilde{X}$ is a Nash equilibrium in the nonatomic game $\tilde{G}$ with $\tilde{p}$ if
${\tilde{x}^{j\ast}_{i}}(\tilde{p})$ $>0,\quad$
$\Rightarrow\tilde{c}_{i}^{j}({\tilde{x}^{\ast}}(\tilde{p}),\tilde{p})\leqslant\tilde{c}_{i}^{j^{\prime}}({\tilde{x}^{\ast}}(\tilde{p}),\tilde{p}),~{}\forall j,j^{\prime}\in\tilde{S}_{i},~{}\forall i\in\tilde{\mathcal{I}}.$
@@
The function $\tilde{\ell}(\cdot)$ is Lipschitz continuous. Moreover, it is strongly monotone: there exists $\rho>0$ such that
~$\left\langle\tilde{\ell}(\tilde{x})-\tilde{\ell}(\tilde{x}^{\prime}),\tilde{x}-\tilde{x}^{\prime}\right\rangle>\rho\|\tilde{x}-\tilde{x}^{\prime}\|^{2},\quad\forall\tilde{x}\neq\tilde{x}^{\prime}\in\tilde{X}.$
Additionally, $\tilde{\Phi}(\tilde{x})$ is twice continuously differentiable, has Lipschitz gradient, and strictly convex.
Under Assumption 2.2, Nash equilibrium $x^{\ast}(p)$ exists, is unique and Lipscthiz continuous for every $p\in\mathbb{R}^{|\tilde{\mathcal{I}}|}$ [Sandholm, 2010, Dafermos, 1988]. Moreover, the socially optimal strategy distribution is also unique.
@@ 2.2 Coupled Strategy and Incentive Update
We consider a coupled dynamics that jointly update agents’ strategies and the incentive mechanism with discrete timesteps $k\in\mathbb{N}$. At step $k$, the strategy profile in the atomic game $G$ (resp. non-atomic game $\tilde{G}$) is $x_{k}=(x_{i,k})_{i\in\mathcal{I}}$ (resp. $\tilde{x}_{k}=(\tilde{x}_{i,k})_{i\in\tilde{\mathcal{I}}}$), where $x_{i,k}$ (resp. $\tilde{x}_{i,k}$) is the strategy of player $i$ (population $i$), and the incentive mechanism is $p_{k}=(p_{i,k})_{i\in\mathcal{I}}$ (resp. $\tilde{p}_{k}=(\tilde{p}_{i,k}^{j})_{j\in S_{i},i\in\tilde{\mathcal{I}}}$). The strategy updates and the incentive updates are presented below:
@@
Some of the examples of commonly studied learning dynamics ($x$-update) and ($\tilde{x}$-update) are delineated below:
@@
    Best response update ([Fudenberg and Levine, 1998, Hofbauer and Sandholm, 2002]): The strategy update incorporates a best response strategy with respect to the strategy and the incentive mechanism in step $k$:
@@
    $\tilde{f}_{i}(\tilde{x}_{k},\tilde{p}_{k})$ $=\underset{\tilde{y}_{i}\in\tilde{X}_{i}}{\arg\min}\ \tilde{y}_{i}^{\top}\tilde{c}_{i}(\tilde{x}_{k},\tilde{p}_{k}).$
@@
    Gradient-based update ([Littlestone and Warmuth, 1994, Laraki and Mertikopoulos, 2013, Sandholm, 2010]): The strategy update incorporates a smoothed strategy update with respect to the strategy and the incentive mechanism in step $k$:
    $f_{i}(x_{k},p_{k})$ $=\underset{y_{i}\in X_{i}}{\arg\min}\ z_{i}(x_{k},p_{k})y_{i}-h(y_{i}),$ (6)
    $\tilde{f}_{i}(\tilde{x}_{k},\tilde{p}_{k})$ $=\underset{\tilde{y}_{i}\in\tilde{X}_{i}}{\arg\min}\ \tilde{c}_{i}(\tilde{x}_{k},\tilde{p}_{k})-\tilde{h}(\tilde{y}_{i}),$
    where $z_{i}(x_{k},p_{k})=x_{k}-\eta\nabla_{x_{i}}c_{i}(x_{k},p_{k})$, $\eta$ is step size, and $h(\cdot),\tilde{h}(\cdot)$ are regularizers. If $h(\cdot)$ is a quadratic function then the update becomes projected gradient descent [Mertikopoulos and Zhou, 2019]. Furthermore, if $\tilde{h}(\cdot)$ is the entropy function then the update becomes a perturbed best-response update[Sandholm, 2010].
@@
$\tilde{e}_{i}^{j}(\tilde{x})$ $=\nabla_{\tilde{x}_{i}^{j}}\tilde{\Phi}(\tilde{x})-\tilde{\ell}_{i}^{j}(\tilde{x}),\quad\forall j\in S_{i},\quad\forall i\in\tilde{\mathcal{I}}.$ (7b)
In (7a), $e_{i}({x})$ is the difference between the marginal social cost, and the marginal cost of player $i$ given ${x}$. Similarly, $\tilde{e}_{i}^{j}(\tilde{x})$ is the difference between the marginal social cost, and the cost experienced by the players in population $i$ who chooses action $j$. We refer $e_{i}({x})$ and $\tilde{e}_{i}(\tilde{x})=(\tilde{e}_{i}^{j}(\tilde{x}))_{j\in S_{i}}$ as the externality of player $i$ and population $i$, respectively, since they capture the difference in impacts of their strategy on the social cost and individual cost.
@@
In summary, the joint evolution of strategy and incentive mechanism $(x_{k},p_{k})_{k=1}^{\infty}$ (resp. $(\tilde{x}_{k},\tilde{p}_{k})_{k=1}^{\infty}$) in the atomic game $G$ (resp. non-atomic game $\tilde{G}$) is governed by the learning dynamics ($x$-update)–($p$-update) (resp. ($\tilde{x}$-update)–($\tilde{p}$-update)). The step-sizes $(\gamma_{k})_{k=1}^{\infty}$ and $(\beta_{k})_{k=1}^{\infty}$ determine the speed of strategy updates and incentive updates.
Remark 2.1.
Our externality-based incentive update contrasts with the gradient-based incentive update considered in the literature [Liu et al., 2021]. Such a gradient-based incentive update typically requires that the equilibrium social cost function $\Phi(x^{\ast}(p))$ (resp. $\tilde{\Phi}(\tilde{x}^{\ast}(\tilde{p}))$) be strongly convex in the incentive mechanism $p$ (resp. $\tilde{p}$). However, this requirement is too restrictive and is not satisfied even in simple games due to the fact that the equilibrium strategy vector is typically non-convex in the incentive mechanism. In Appendix A, we provide an example of a two-link routing game, where the equilibrium social cost function is non-convex in the incentive mechanism, and the gradient-based learning dynamics do not converge.
Remark 2.2.
The incentive update is based on only limited information about the players’ behavior. Specifically, it relies solely on the gradients of the social cost and those of the players’, evaluated at the players’ current strategy profile. Note that it is agnostic to the strategy update dynamics (i.e., ($x$-update), and ($\tilde{x}$-update)) employed by players.
@@ 3 General results
In Sec. 3.1, we characterize the set of fixed points of the updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update), and show that any fixed point corresponds to a socially optimal incentive mechanism such that the induced Nash equilibrium strategy profile minimizes the social cost. In Sec. 3.2, we provide a set of sufficient conditions that guarantee the convergence of strategy and incentive updates. Under these conditions, our adaptive incentive mechanism eventually induces a socially optimal outcome.
@@ 3.1 Fixed point analysis
Our first result characterizes conditions under which the fixed point set $P^{\dagger}$ (resp. $\tilde{{P}}^{\dagger}$) is non-empty and singleton in $G$ (resp. $\tilde{G}$). Moreover, given any fixed point incentive parameter $p^{\dagger}\in P^{\dagger}$ and $\tilde{p}^{\dagger}\in\tilde{{P}}^{\dagger}$, the corresponding Nash equilibrium is socially optimal.
@@
Proof.
First, we show that $P^{\dagger}$ is non-empty. That is, there exists $p^{\dagger}$ such that $e(x^{\ast}(p^{\dagger}))=p^{\dagger}$. Define a function $\theta(p)=e(x^{\ast}(p))$. We note that Assumption 2.1 guarantees that $\theta$ is well defined. Thus the problem is reduced to showing existence of fixed point of $\theta(\cdot)$, for which we use the Schauder fixed point theorem [Smart, 1980].
We note that Assumption 2.1 ensures that $\theta(p)$ is a continuous function. Now, define $K\coloneqq\{\theta(p):p\in\mathbb{R}^{|\mathcal{I}|}\}\subseteq\mathbb{R}^{|\mathcal{I}|}$. We claim that the set $K$ is compact. Indeed, this follows by two observations. First, the externality function $e(\cdot)$ is continuous. Second, the range of the function $x^{\ast}(\cdot)$ is $X$ which is a compact set. These two observations ensure that $\theta(p)=e(x^{\ast}(p))$ is a bounded function. Let $\tilde{K}\coloneqq\textsf{conv}(K)$ be the convex hull of $K$, which in turn is also a compact set. Let’s denote the restriction of function $\theta$ on the set $\tilde{K}$ as $\theta_{|\tilde{K}}:\tilde{K}\longrightarrow\tilde{K}$ where $\theta_{|\tilde{K}}(p)=\theta(p)$ for all $p\in\tilde{K}$. We note that $\theta_{|\tilde{K}}$ is a continuous function from a convex compact set to itself and therefore, the Schauder fixed point theorem ensures that there exists $p^{\dagger}\in\tilde{K}$ such that $p^{\dagger}=\theta_{|\tilde{K}}(p^{\dagger})=\theta(p^{\dagger})$. This concludes the proof of existence of $p^{\dagger}$. A similar line of arguments as above can be replicated for non-atomic game $\tilde{G}$ to show that $\tilde{P}^{\dagger}$ is non-empty. We omit the details here for the sake of brevity.
Next, we show that the incentive $p^{\dagger}$ aligns the Nash equilibrium with social optimality (i.e. for any $p^{\dagger}\in P^{\dagger}$, $x^{\ast}(p^{\dagger})={x}^{\dagger}$). Fix $p^{\dagger}\in P^{\dagger}$. For every $i\in\mathcal{I}$ we have $p^{\dagger}_{i}=e_{i}(x^{\ast}(p^{\dagger}))$. This implies $\nabla_{{x}_{i}}\ell_{i}(x^{\ast}(p^{\dagger}))+p^{\dagger}_{i}=\nabla_{{x}_{i}}\Phi(x^{\ast}(p^{\dagger}))$ for every $i\in\mathcal{I}$. This implies
@@
where $J(x,p)$ is the game Jacobian defined as $J_{i}(x,p)=\nabla_{{x}_{i}}\ell_{i}(x)+p_{i}$ for every $i\in\mathcal{I}$. Next, under Assumption 2.1, the variational inequality characterization of Nash equilibrium [Facchinei and Pang, 2007], we know that $x^{\ast}(p^{\dagger})$ is a Nash equilibrium if and only if
@@
From (10) and (11) observe that
@@
Comparing (12) with (13), we note that $x^{\ast}(p^{\dagger})$ is the minimizer of social cost function $\Phi$. This implies $x^{\ast}(p^{\dagger})={x}^{\dagger}$ as ${x}^{\dagger}$ is the unique minimizer of social cost function $\Phi$ by Assumption 2.1.
Similarly for non-atomic game $\tilde{G}$, we show that the incentive $\tilde{p}^{\dagger}$ aligns the Nash equilibrium with social optimality (i.e. for any $\tilde{p}^{\dagger}\in\tilde{{P}}^{\dagger}$, ${\tilde{x}^{\ast}}(\tilde{p}^{\dagger})={\tilde{x}}^{\dagger}$). Fix $\tilde{p}^{\dagger}\in\tilde{{P}}^{\dagger}$. For every $j\in\tilde{S}_{i}$ and $i\in\tilde{\mathcal{I}}$, it holds that $\tilde{p}^{j\dagger}_{i}=\tilde{e}_{i}^{j}({\tilde{x}^{\ast}}(\tilde{p}^{\dagger}))$. This implies that
@@
Under Assumption 2.2, the variational inequality characterization of Nash equilibrium for non-atomic games [Sandholm, 2010], we know that ${\tilde{x}^{\ast}}(\tilde{p}^{\dagger})$ is a Nash equilibrium if and only if
$\langle\tilde{c}({\tilde{x}^{\ast}}(\tilde{p}^{\dagger}),\tilde{p}^{\dagger}),\tilde{x}-{\tilde{x}^{\ast}}(\tilde{p}^{\dagger})\rangle\geqslant 0\quad\forall\ \tilde{x}\in\tilde{X}.$ (15)

### v3

@@ 2.1.2 Non-atomic Games
Consider a game $\tilde{G}$ with a finite set of player populations $\tilde{\mathcal{I}}$. Each population $i\in\tilde{\mathcal{I}}$ is comprised of a continuum set of (infinitesimal) players with mass $\tilde{M}_{i}>0$. Every (infinitesimal) player in any population can choose an action in a finite set $\tilde{S}_{i}$. The strategy distribution of population $i\in\tilde{\mathcal{I}}$ is $\tilde{x}_{i}=(\tilde{x}_{i}^{j})_{j\in\tilde{S}_{i}}$, where $\tilde{x}_{i}^{j}$ is the mass of players in population $i$ who choose action $j\in\tilde{S}_{i}$. Then, the set of all strategy distributions of population $i$ is $\tilde{X}_{i}=\left\{\tilde{x}_{i}\in\mathbb{R}^{|\tilde{S}_{i}|}|\sum_{j\in\tilde{S}_{i}}\tilde{x}_{i}^{j}=\tilde{M}_{i},~{}\tilde{x}_{i}^{j}\geqslant 0,\forall j\in\tilde{S}_{i}\right\}$. The strategy distribution of all populations is given by $\tilde{x}=(\tilde{x}_{i})_{i\in\tilde{\mathcal{I}}}\in\tilde{X}=\prod_{i\in\tilde{\mathcal{I}}}\tilde{X}_{i}$. We define $\tilde{S}=\prod_{i\in\tilde{\mathcal{I}}}\tilde{S}_{i}$. Given a strategy distribution $\tilde{x}\in\tilde{X}$, the cost of players in population $i\in\tilde{\mathcal{I}}$ for choosing action $j\in\tilde{S}_{i}$ is ~$\tilde{\ell}_{i}^{j}(\tilde{x})$. We denote ~$\tilde{\ell}_{i}(\tilde{x})=(\tilde{\ell}_{i}^{j}(\tilde{x}))_{j\in\tilde{S}_{i}}$ as the vector of costs for each $i\in\tilde{\mathcal{I}}$.
A system operator designs incentives by setting a payment $\tilde{p}_{i}^{j}$ for players in population $i$ who choose action $j\in\tilde{S}_{i}$. Consequently, given the incentive mechanism $\tilde{p}=({\tilde{p}_{i}^{j}})_{j\in S_{i},i\in\tilde{\mathcal{I}}}$, the total cost experienced by any player in population $i\in\tilde{\mathcal{I}}$ who chooses action $j\in\tilde{S}_{i}$ is
@@
A strategy distribution ${\tilde{x}^{\ast}}(\tilde{p})\in\tilde{X}$ is a Nash equilibrium in the non-atomic game $\tilde{G}$ with $\tilde{p}$ if
${\tilde{x}^{j\ast}_{i}}(\tilde{p})$ $>0,\quad\Rightarrow\tilde{c}_{i}^{j}({\tilde{x}^{\ast}}(\tilde{p}),\tilde{p})\leqslant\tilde{c}_{i}^{j^{\prime}}({\tilde{x}^{\ast}}(\tilde{p}),\tilde{p}),$ (3)
$\hskip 56.9055pt~{}\forall j,j^{\prime}\in\tilde{S}_{i},~{}\forall i\in\tilde{\mathcal{I}}.$
@@
For any $p\in\mathbb{R}^{|\tilde{\mathcal{I}}|}$, the Nash equilibrium $\tilde{x}^{\ast}(\tilde{p})$ is unique and Lipschitz continuous in $\tilde{p}$. Moreover, $\tilde{\Phi}(\tilde{x})$ is continuously differentiable and strictly convex.
Assumption 2.2 is widely adopted in the literature on incentive design for non-atomic games (e.g., [Liu et al., 2022, Ochoa and Poveda, 2022, Mojica-Nava et al., 2022]), either directly or through other conditions that guarantee this [^2] Uniqueness and Lipschitz continuity of $x^{\ast}(p)$ hold if $\tilde{\ell}(\cdot)$ is Lipschitz continuous and strongly monotone [Sandholm, 2010]. That is, there exists $\rho>0$ such that ~$\langle\tilde{\ell}(\tilde{x})-\tilde{\ell}(\tilde{x}^{\prime}),\tilde{x}-\tilde{x}^{\prime}\rangle\geqslant\rho\|\tilde{x}-\tilde{x}^{\prime}\|^{2}$ for every $\tilde{x}\neq\tilde{x}^{\prime}\in\tilde{X}.$
@@ 2.2 Coupled Strategy and Incentive Update
We consider a coupled dynamical system that jointly updates players’ strategies and the incentive mechanism with discrete time-steps $k\in\mathbb{N}$. At step $k$, the strategy profile in the atomic game $G$ (resp. non-atomic game $\tilde{G}$) is $x_{k}=(x_{i,k})_{i\in\mathcal{I}}$ (resp. $\tilde{x}_{k}=(\tilde{x}_{i,k})_{i\in\tilde{\mathcal{I}}}$), where $x_{i,k}$ (resp. $\tilde{x}_{i,k}$) is the strategy of player $i$ (population $i$), and the incentive mechanism is $p_{k}=(p_{i,k})_{i\in\mathcal{I}}$ (resp. $\tilde{p}_{k}=(\tilde{p}_{i,k}^{j})_{j\in S_{i},i\in\tilde{\mathcal{I}}}$). The strategy updates and the incentive updates are presented below:
@@
We shall impose additional assumptions on $f(\cdot)$ and $\tilde{f}(\cdot)$ when studying the convergence of strategy and incentive updates in the next section. Some examples of commonly studied learning dynamics ($x$-update) and ($\tilde{x}$-update) include:
@@
    Best response update ([Fudenberg and Levine, 1998, Sandholm, 2010]): The strategy update incorporates a best response strategy with respect to the strategy and the incentive mechanism in step $k$:
@@
    $\tilde{f}_{i}(\tilde{x}_{k},\tilde{p}_{k})$ $=\underset{\tilde{y}_{i}\in\tilde{X}_{i}}{\arg\min}\ \tilde{y}_{i}^{\top}\tilde{c}_{i}(\tilde{x}_{k},\tilde{p}_{k}),$
    where the first equation is the best response update in atomic games [Fudenberg and Levine, 1998], and the second is the best response update in non-atomic games [Sandholm, 2010].
@@
    Gradient-based update ([Littlestone and Warmuth, 1994, Laraki and Mertikopoulos, 2013, Sandholm, 2010]): Gradient-based strategy update commonly studied in literature takes the following form:
    $f_{i}(x_{k},p_{k})$ $=\underset{y_{i}\in X_{i}}{\arg\max}\ z_{i}(x_{k},p_{k})y_{i}-h(y_{i}),$ (7)
    $\tilde{f}_{i}(\tilde{x}_{k},\tilde{p}_{k})$ $=\underset{\tilde{y}_{i}\in\tilde{X}_{i}}{\arg\max}\ {{}\tilde{y}_{i}^{\top}}\tilde{c}_{i}(\tilde{x}_{k},\tilde{p}_{k})-\tilde{h}(\tilde{y}_{i}),$
    where $z_{i}(x_{k},p_{k})=x_{k}-\eta\nabla_{x_{i}}c_{i}(x_{k},p_{k})$, $\eta$ is step size, and $h(\cdot),\tilde{h}(\cdot)$ are regularizers. If $h(\cdot)$ is a quadratic function, then the update becomes projected gradient descent [Mertikopoulos and Zhou, 2019]. Furthermore, if $\tilde{h}(\cdot)$ is the entropy function, then the update becomes a perturbed best-response update[Sandholm, 2010].
@@
$\tilde{e}_{i}^{j}(\tilde{x})$ $=\nabla_{\tilde{x}_{i}^{j}}\tilde{\Phi}(\tilde{x})-\tilde{\ell}_{i}^{j}(\tilde{x}),\quad\forall j\in\tilde{S}_{i},\quad\forall i\in\tilde{\mathcal{I}}.$ (8b)
In (8a), $e_{i}({x})$ represents the difference between the marginal social cost and the marginal cost of player $i$ given ${x}$. Similarly, $\tilde{e}_{i}^{j}(\tilde{x})$ denotes the difference between the marginal social cost and the cost experienced by players in population $i$ who choose action $j$. We refer to $e_{i}({x})$ and $\tilde{e}_{i}(\tilde{x})=(\tilde{e}_{i}^{j}(\tilde{x}))_{j\in S_{i}}$ as the externalities of players $i$ and population $i$, respectively, since they capture the difference in the impact of their strategies on the social cost and individual cost.
@@
In summary, the joint evolution of strategy and incentive mechanism $(x_{k},p_{k})_{k=1}^{\infty}$ (resp. $(\tilde{x}_{k},\tilde{p}_{k})_{k=1}^{\infty}$) in the atomic game $G$ (resp. non-atomic game $\tilde{G}$) is governed by the learning dynamics ($x$-update)–($p$-update) (resp. ($\tilde{x}$-update)–($\tilde{p}$-update)). The step-sizes $(\gamma_{k})_{k=1}^{\infty}$ and $(\beta_{k})_{k=1}^{\infty}$ determine the speed of strategy updates and incentive updates, respectively.
Information environment of incentive update. The incentive updates in ($p$-update) and ($\tilde{p}$-update) are based on the externalities created by players’ strategies. In the absence of additional problem structure, computing externality requires oracle access to the gradient of players’ costs (first-order information) in atomic games (cf. (8a)) or the players’ cost functions (zeroth-order information) in non-atomic games (cf. (8b)), both evaluated at the current strategy profile. This information requirement is less demanding compared to the gradient-based incentive updates adopted in previous literature [Mojica-Nava et al., 2022, Chen et al., 2023, Poveda et al., 2017, Ochoa and Poveda, 2022, Liu et al., 2022, Li et al., 2023, Alpcan and Pavel, 2009, Alpcan et al., 2009], where estimating the gradient of the social cost function with respect to the incentive vector often requires the knowledge of the game Jacobian (i.e., second-order information) 333 For instance, the gradient based incentive update of atomic games studied in [Liu et al., 2022] takes the following form $p_{k+1}=p_{k}-\beta_{k}\nabla_{p}x^{\ast}(p_{k})^{\top}\nabla_{x}\Phi(x^{\ast}(p_{k})),$ which is a gradient descent update on the function $\Phi(x^{\ast}(p))$. The authors estimate $\nabla_{p}x^{\ast}(p_{k})^{\top}$ with $-\nabla_{p}J(x_{k};p_{k})^{\top}(\nabla_{x}J(x_{k};p_{k}))^{-1}$, where $J(x;p)=(\partial c_{i}(x,p)/\partial x_{i})_{i\in\mathcal{I}}$ is the game Jacobian. Therefore, these updates require second order information about the cost function of players. Meanwhile, our approach of externality based pricing only requires first-order information about the cost function of players. [Liu et al., 2022, Li et al., 2023], or knowledge of the gradient of equilibrium strategy with respect to incentive (i.e., $\nabla_{p}x^{\ast}(p)$) [Alpcan et al., 2009, Alpcan and Pavel, 2009]. Furthermore, our incentive updates do not require knowledge of players’ entire cost function, and are agnostic to the specific strategy update dynamics (i.e., ($x$-update) and ($\tilde{x}$-update)) employed by the players.
In many settings, leveraging the structure of the underlying problem enables the social planner to compute externalities with less information. For instance, in non-atomic routing games (see Section 4.2), the social planner can compute externality using only the travel time costs of edges (road segments in the network) instead of the cost of each path taken by each population given their origin-destination pair. Additionally, in energy system applications (e.g., [Li et al., 2024]), player’s cost function $\ell_{i}(x)=g_{i}(x_{i})$ often only depends on their own energy consumption $x_{i}$, and the social cost function $\Phi(x)=r(x)+\sum_{i\in\mathcal{I}}g_{i}(x_{i})$ is modeled as the sum of the public cost $r(x)$ that depends on the joint action $x$ and the cost of individual players. In this case, the externality for any player $i$ depends only on the gradient of the public cost function $r(x)$ and not on the private cost of players.
$e_{i}(x)=\frac{\partial\Phi(x)}{\partial x_{i}}-\frac{\partial\ell_{i}(x)}{\partial x_{i}}=\frac{\partial r(x)}{\partial x_{i}}.$
@@ 3 General results
In Section 3.1, we characterize the set of fixed points of the updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update), and show that any fixed point corresponds to a socially optimal incentive mechanism such that the induced Nash equilibrium strategy profile minimizes the social cost. In Section 3.2, we provide a set of sufficient conditions that guarantee (local and global) convergence of incentive updates. Under these conditions, our adaptive incentive mechanism eventually induces a socially optimal outcome.
@@ 3.1 Fixed point analysis
Our first result characterizes conditions under which the fixed point set $P^{\dagger}$ (resp. $\tilde{{P}}^{\dagger}$) is non-empty and singleton in $G$ (resp. $\tilde{G}$). Moreover, given any fixed point incentive mechanism $p^{\dagger}\in P^{\dagger}$ and $\tilde{p}^{\dagger}\in\tilde{{P}}^{\dagger}$, the corresponding Nash equilibrium is socially optimal.
@@
Advantage of externality-based incentive updates. Proposition 3.1 demonstrates that the externality-based incentive updates ($p$-update) and ($\tilde{p}$-update) ensure that any fixed point must achieve social optimality. In contrast, the gradient-based incentive update, commonly considered in the literature (e.g., [Liu et al., 2022, Li et al., 2023, Mojica-Nava et al., 2022, Alpcan et al., 2009, Alpcan and Pavel, 2009]), does not guarantee that its fixed point corresponds to a socially optimal incentive mechanism. Typically, these works impose additional assumptions, such as the equilibrium social cost function $\Phi(x^{\ast}(p))$ (resp. $\tilde{\Phi}(\tilde{x}^{\ast}(\tilde{p}))$) being strongly convex in the incentive mechanism $p$ (resp. $\tilde{p}$) [Liu et al., 2022, Li et al., 2023, Mojica-Nava et al., 2022], or that the gradient of the equilibrium strategy, $\nabla_{{p}}{x}^{\ast}({p})$ (resp. $\nabla_{\tilde{p}}\tilde{x}^{\ast}(\tilde{p})$), with respect to the incentive mechanism $p$ (resp. $\tilde{p}$) is non-singular [Alpcan et al., 2009, Alpcan and Pavel, 2009], to ensure that the fixed points of the gradient-based update achieve the socially optimal outcome. In fact, in Appendix A, we provide an example of a two-link non-atomic routing game, where these assumptions are not satisfied and nearly all fixed points of the gradient-based incentive update fail to achieve the socially optimal outcome. Consequently, the gradient-based incentive update can lead to inefficient outcomes. In contrast, our externality-based incentive update has a unique fixed point that always induces the socially optimal outcome.
Proof of Proposition 3.1. First, we show that $P^{\dagger}$ is non-empty, i.e., there exists $p^{\dagger}$ such that $e(x^{\ast}(p^{\dagger}))=p^{\dagger}$. Define the function $\theta(p)=e(x^{\ast}(p))$. By Assumption 2.1, $\theta$ is well-defined. Thus, the problem reduces to proving the existence of a solution to $p=\theta(p)$.
We note that Assumption 2.1 ensures that $\theta(p)$ is a continuous function. Now, define $K\coloneqq\{\theta(p):p\in\mathbb{R}^{|\mathcal{I}|}\}\subseteq\mathbb{R}^{|\mathcal{I}|}$. We claim that the set $K$ is compact. Indeed, this follows from two observations. First, the externality function $e(\cdot)$ is continuous. Second, the range of the function $x^{\ast}(\cdot)$ is $X$, which is a compact set. These two observations ensure that $\theta(p)=e(x^{\ast}(p))$ is a bounded function. Let $\tilde{K}\coloneqq\textsf{conv}(K)$ be the convex hull of $K$, which in turn is also a compact set. Let’s denote the restriction of function $\theta$ on the set $\tilde{K}$ as $\theta_{|\tilde{K}}:\tilde{K}\longrightarrow\tilde{K}$ where $\theta_{|\tilde{K}}(p)=\theta(p)$ for all $p\in\tilde{K}$. We note that $\theta_{|\tilde{K}}$ is a continuous function from a convex compact set to itself and therefore, the Schauder fixed point theorem ensures that there exists $p^{\dagger}\in\tilde{K}$ such that $p^{\dagger}=\theta_{|\tilde{K}}(p^{\dagger})=\theta(p^{\dagger})$ [Smart, 1980]. This concludes the proof of the existence of $p^{\dagger}$. Analogous argument applies for the non-atomic game $\tilde{G}$ to show that $\tilde{P}^{\dagger}$ is non-empty.
Next, we show that the incentive $p^{\dagger}$ aligns the Nash equilibrium with socially optimal strategy (i.e. for any $p^{\dagger}\in P^{\dagger}$, $x^{\ast}(p^{\dagger})={x}^{\dagger}$). For any $p^{\dagger}\in P^{\dagger}$ and any $i\in\mathcal{I}$, it holds that $p^{\dagger}_{i}=e_{i}(x^{\ast}(p^{\dagger}))$. This implies that $\nabla_{{x}_{i}}\ell_{i}(x^{\ast}(p^{\dagger}))+p^{\dagger}_{i}=\nabla_{{x}_{i}}\Phi(x^{\ast}(p^{\dagger}))$ for every $i\in\mathcal{I}$, and thus
@@
where $J(x,p)$ is the game Jacobian defined as $J_{i}(x,p)=\nabla_{{x}_{i}}\ell_{i}(x)+p_{i}$ for every $i\in\mathcal{I}$. From Assumption 2.1 and the first order necessary condition for Nash equilibrium [Facchinei and Pang, 2007], we know that the Nash equilibrium $x^{\ast}(p^{\dagger})$ must satisfy
@@
From (11) and (12), we observe that
@@
Comparing (13) with (14), we note that $x^{\ast}(p^{\dagger})$ is the minimizer of social cost function $\Phi$. This implies that $x^{\ast}(p^{\dagger})={x}^{\dagger}$, since ${x}^{\dagger}$ is the unique minimizer of the social cost function $\Phi$ under Assumption 2.1.
Similarly, for non-atomic game $\tilde{G}$, we show that the incentive $\tilde{p}^{\dagger}$ aligns the Nash equilibrium with social optimality. Fix $\tilde{p}^{\dagger}\in\tilde{{P}}^{\dagger}$. For every $j\in\tilde{S}_{i}$ and $i\in\tilde{\mathcal{I}}$, it holds that $\tilde{p}^{j\dagger}_{i}=\tilde{e}_{i}^{j}({\tilde{x}^{\ast}}(\tilde{p}^{\dagger}))$. Consequently,
@@
Under Assumption 2.2, ${\tilde{x}^{\ast}}(\tilde{p}^{\dagger})$ is a Nash equilibrium only if
$\langle\tilde{c}({\tilde{x}^{\ast}}(\tilde{p}^{\dagger}),\tilde{p}^{\dagger}),\tilde{x}-{\tilde{x}^{\ast}}(\tilde{p}^{\dagger})\rangle\geqslant 0,\quad\forall\ \tilde{x}\in\tilde{X}.$ (16)